OpenStudy (anonymous):

Helppp please i will give medals Simplify completely: 14x^5y^4+21x^3y^2 ------------------- 7x^3y answers 2x8y5 + 3x6y3 2x2y4 + 3xy 2x2y3 + 3y 2x2y3 + 3y2

OpenStudy (nurali):

\[\frac{ 14x^5y^4+21x^3y^2 }{ 7x^3y }\] \[\frac{ 14x^5y^4 }{ 7x^3y }+\frac{ 21x^3y^2 }{ 7x^3y }\] \[ 2x^{5-3} y ^{4-1}+3x ^{3-3} y ^{2-1}\] \[2x^2y^3+3y\]

OpenStudy (anonymous):

Create a unique example of dividing a polynomial by a monomial and provide the simplified form. Explain, in complete sentences, the two ways used to simplify this expression and how you would check your quotient for accuracy.